OpenStudy (anonymous):

(-Π /4)sin(Π t/4)=0, solve for t

OpenStudy (anonymous):

\[\frac{ \pi t}{4} = n \pi \]

OpenStudy (anonymous):

\[t=4n\]

OpenStudy (anonymous):

for integers n

OpenStudy (anonymous):

how did you get Πt/4=nΠ